OpenStudy (anonymous):

given the equation 6x-y=98,find x if y =-2

OpenStudy (anonymous):

\[6x = 98 + y\] do you agree?

HanAkoSolo (jamierox4ev3r):

also, if you want another process, plyg in the y-value taht is given :) 6x-y=98 = 6x+2=98 once the equation is simplified to 6x+2=98, you can solve for x 6x+2=98 -2 -2 -------- 6x=96 /6 /6 ------ x=16

OpenStudy (uri):

Uri is here :D So, Put -2 instead of y 6x-(-2)=98 6x+2=98 Subract 2 both sides 6x+2=98 -2 -2 ======== 6x=96 Divide 96 with 6 x=96/6 x=16 ans @Jordan01